# FedAdamW: A Communication-Efficient Optimizer with Convergence and Generalization Guarantees for Federated Large Models

### v1

@@ Appendix C Appendix C: Experimental Setup
  • CIFAR-100 (Krizhevsky, Hinton et al. 2009): Contains 100 classes with 600 color images per class at a resolution of $32\times 32$. It is a standard benchmark for evaluating federated image classification methods.
@@
We adopt ResNet-18 as the backbone model. To better adapt it to CIFAR-100, we modify its architecture following standard practices (He et al. 2016):
@@
We construct a lightweight Vision Transformer model, ViT-Tiny, specifically tailored for federated learning on the CIFAR-100 dataset. The design is based on the standard ViT architecture (Dosovitskiy et al. 2020), with modifications to accommodate the small input size and limited data per client.
@@
To demonstrate the effectiveness of our method on large-scale vision models, we conduct fine-tuning experiments using Swin Transformer-Tiny and ViT-Base on Tiny ImageNet and CIFAR-100. For both models, we initialize from official ImageNet-22K pre-trained weights (Liu et al. 2021; Dosovitskiy et al. 2020) to ensure consistency across methods.
@@ D.1 Motivation and Background
Figure 8: The near-block-diagonal Hessian structure of neural nets in FL. (a) is the Hessian of an MLP reported in (Collobert 2004). (b–d) are the Hessians of a 1-hidden-layer MLP on CIFAR-100. The near-block-diagonal structure maintains throughout training in FL.
@@ D.2 Practical Partitioning Strategy
  • Class 3: value. For value, the structure of the Hessian seems less clear. It appears to have 16 diagonal blocks (matching the number of output neurons), but the pattern is less obvious. This structure is significantly different from query and key, despite all using four heads. Additionally, the Hessian entries of value are about $10^{6}$ larger than those of query and key [^1] This might be one source of the heterogeneity of Hessian eigenvalues as reported by Zhang et al. (2024a). One possible reason is that value is positioned outside the softmax operator in the self-attention design, unlike query and key.

### v2

@@ Appendix C Appendix C: Experimental Setup
  • CIFAR-100 (krizhevsky2009learning): Contains 100 classes with 600 color images per class at a resolution of $32\times 32$. It is a standard benchmark for evaluating federated image classification methods.
@@
We adopt ResNet-18 as the backbone model. To better adapt it to CIFAR-100, we modify its architecture following standard practices (he2016deep):
@@
We construct a lightweight Vision Transformer model, ViT-Tiny, specifically tailored for federated learning on the CIFAR-100 dataset. The design is based on the standard ViT architecture (dosovitskiy2020image), with modifications to accommodate the small input size and limited data per client.
@@
To demonstrate the effectiveness of our method on large-scale vision models, we conduct fine-tuning experiments using Swin Transformer-Tiny and ViT-Base on Tiny ImageNet and CIFAR-100. For both models, we initialize from official ImageNet-22K pre-trained weights (liu2021swin; dosovitskiy2020image) to ensure consistency across methods.
@@ D.1 Motivation and Background
Figure 8: The near-block-diagonal Hessian structure of neural nets in FL. (a) is the Hessian of an MLP reported in (collobert2004large). (b–d) are the Hessians of a 1-hidden-layer MLP on CIFAR-100. The near-block-diagonal structure maintains throughout training in FL.
@@ D.2 Practical Partitioning Strategy
  • Class 3: value. For value, the structure of the Hessian seems less clear. It appears to have 16 diagonal blocks (matching the number of output neurons), but the pattern is less obvious. This structure is significantly different from query and key, despite all using four heads. Additionally, the Hessian entries of value are about $10^{6}$ larger than those of query and key [^1] This might be one source of the heterogeneity of Hessian eigenvalues as reported by zhang2024transformers. One possible reason is that value is positioned outside the softmax operator in the self-attention design, unlike query and key.